Two trucks travel from city A to city B. Truck A travels 10mphs faster then truck B. Truck A takes 4hrs and truck B travels 5hrs to make the same trip. How far apart are the 2 cities?
Answer by mananth(16949)

distance A to B = x
speed of Truck B = y
speed of truck A = y+10
Truck B
x/y = 5
x=5y
Truck A
x/(y+10)= 4
x=4(y+10)
x=4y+40
equate x
5y=4y+40
y=40 mph, speed of truck B
Truck A speed = 50 mph
plug y in x=5y
x= 5*40 =200 miles
m.ananth@hotmail.ca